# Training Task Reasoning LLM Agents for Multi-turn Task Planning via Single-turn Reinforcement Learning

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in knowledge acquisition, reasoning, and tool use, making them promising candidates for autonomous agent applications. However, training LLM agents for complex multi-turn task planning faces significant challenges, including sparse episode-wise rewards, credit assignment across long horizons, and the computational overhead of reinforcement learning in multi-turn interaction settings. To this end, this paper introduces a novel approach that transforms multi-turn task planning into single-turn task reasoning problems, enabling efficient policy optimization through Group Relative Policy Optimization (GRPO) with dense and verifiable reward from expert trajectories. Our theoretical analysis shows that GRPO improvement on single-turn task reasoning results in higher multi-turn success probability under the minimal turns, as well as the generalization to subtasks with shorter horizons. Experimental evaluation on the complex task planning benchmark demonstrates that our 1.5B parameter model trained with single-turn GRPO achieves superior performance compared to larger baseline models up to 14B parameters, with success rates of 70% for long-horizon planning tasks with over 30 steps. We also theoretically and empirically validate the strong cross-task generalizability that the models trained on complex tasks can lead to the successful completion of all simpler subtasks.
@@ 1 Introduction
  • We provide a theoretical analysis bridging single-turn task reasoning and multi-turn task completion, proving that single-turn GRPO improves the success probability in multi-turn planning scenarios.
@@ 2.2 Expert Trajectory with Unique Optimality
On top of Definition˜2.1, we further define the expert trajectory generated by the expert policy with the unique optimality of minimal trajectory length.
@@
Even though expert trajectory $\tau^{GT}(T^{GT})$ is uniquely optimal with minimal length of turns $T^{GT}$, the induced expert policy $\pi^{GT}$ is not necessarily unique or deterministic, as there is no state-action constraint beyond expert trajectories $\tau^{GT}(T^{GT})$.
@@ 3 Methodology
We first give the theoretical analysis of policy learning for single-turn MDP $\mathcal{M}_{S}$ based on expert trajectories. Then we define the success probability with the minimal turns for task planning MDP $\mathcal{M}$ and show that policy improvement from $\mathcal{M}_{S}$ can lead to higher success probability in $\mathcal{M}$. Finally, we show that the learned policy can generalize well to unseen task planning with fewer completion turns.
@@ 3.1.2 GRPO Single-Turn Optimality and Success Amplification
Proof can be found in Appendix D of [19]. ∎
@@ 3.2 Success Probability for Multi-Turn Task Planning
Please note that the recursion equation in Theorem˜3.2 is not the same as Bellman recursion in value function iteration [20], and the key difference lies in that the multiplication of the binary single-turn reward $r_{\pi^{GT}}(s_{t},a)$ is involved in the recursion of expectation, instead of addition. This recursion bridges the final multi-turn success in Equation˜8 with single-turn reward improvements at each step in Equation˜7, which is crucial to the following analysis on how single-turn GRPO improves the multi-turn task planning.
3.3 GRPO Improvement of Multi-Turn Success Probability
We now analyze how GRPO improvements on single-turn task reasoning in $\mathcal{M}_{S}$ will lead to performance gains in the multi-turn task planning MDP $\mathcal{M}$. First, we assume the GRPO-optimized policy for single-turn task reasoning based on expert trajectories can generalize to all states.
Assumption 3.1 (GRPO improvement on all states).
The GRPO improvement on expert trajectory states in Equation˜7 also holds for all states $s\in\mathcal{S}$, i.e., $\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}[r_{\pi^{GT}}(s,a)]\geq\mathbb{E}_{a\sim\pi^{\text{ref}}(\cdot|s)}[r_{\pi^{GT}}(s,a)],$ where $\pi^{*}$ is the GRPO-optimized policy from Corollary˜3.1.
The assumption above is a slight relaxation of Corollary˜3.1, which empirically holds when the policy model is fine-tuned from large language models (LLMs) with strong generalizability [21, 22]. We then present how the GRPO will improve the multi-turn task planning with a higher success probability with the fewest turns.
Theorem 3.3 (GRPO improves success probability).
@@ 3.3 GRPO Improvement of Multi-Turn Success Probability
$P_{t}^{\pi^{*}}(s_{t})\geq P_{t}^{\pi^{\text{ref}}}(s_{t})\quad\forall s_{t}\in\mathcal{S}.$ (16)
@@
By the existence and uniqueness of the expert trajectory in Definition˜2.2, denote the expert trajectory starting from $s_{t}$ as $\tau^{GT}=[(s_{t},a_{t}),\dots,(s_{T^{GT}},a_{T^{GT}})]$ with minimal turns of task completion $T^{*}(s_{t})=T^{GT}-t$ in Definition˜3.1. We prove by backward induction on the length of remaining turns to task completion. In the final case when $t=T^{GT}$, we have $P_{t}^{\pi^{*}}(s_{t})=P_{t}^{\pi^{\text{ref}}}(s_{t})=1$ since the task is already completed. Now for the inductive case at step $t+1,t+1\leq T^{GT}$, we assume $P_{t+1}^{\pi^{*}}(s^{\prime})\geq P_{t+1}^{\pi^{\text{ref}}}(s^{\prime})$ for all states $s^{\prime}$. Then for any state $s_{t}$ at step $t$, it holds that
@@
$\geq\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}[r_{\pi^{GT}}(s,a)\cdot P_{t+1}^{\pi^{\text{ref}}}(f(s,a))]$ (18)
$\geq\mathbb{E}_{a\sim\pi^{\text{ref}}(\cdot|s)}[r_{\pi^{GT}}(s,a)\cdot P_{t+1}^{\pi^{\text{ref}}}(f(s,a))]$ (19)
$=P_{t}^{\pi^{\text{ref}}}(s),$ (20)
where Equations˜17 and 20 are from Theorem˜3.2, Equation˜18 is based on the inductive hypothesis at step $t+1$ and Equation˜19 is based on Assumption 3.1. ∎
@@
For any subtask defined by completion step $k^{*}<T^{GT}$ in Definition˜3.3, the GRPO-optimized policy $\pi^{*}$ achieves better performance on the subtask than the reference policy $\pi^{\text{ref}}$ for any given state $s_{t}$ at step $t\leq k^{*}$, i.e., $P_{t}^{\text{sub},\pi^{*}}(s_{t})\geq P_{t}^{\text{sub},\pi^{\text{ref}}}(s_{t})$.
@@ 4 Experiments
In the experiments, we aim to answer the following two questions empirically: 1) How is the performance of multi-turn task planning after GRPO over single-turn task reasoning? 2) Can the LLM agents after GRPO generalize to unseen planning tasks? We answer the first question in Section˜4.2 and the second one in Section˜4.3, which correspond to Theorem˜3.3 and Corollary˜3.2, respectively. Prior to the answers, we first introduce the experimental setup of environment, model training and evaluation metrics.
@@ 4.1 Experimental Setup
We evaluate our approach on the Robotouille benchmark [23], a challenging multi-turn task planning environment designed for testing LLM agents’ reasoning capabilities in cooking scenarios. Robotouille provides a deterministic long-horizon complex environment where agents must coordinate multiple actions across different kitchen stations to complete recipes. We focus on four progressively challenging cooking tasks with different maximum steps: (1) Cheese Sandwich - at most 10 steps; (2) Burger - at most 10 steps; (3) Cheese Burger - at most 15 steps; and (4) Double Cheese Burger - at most 23 steps. These tasks were selected to provide increasing complexity in both planning horizon and reasoning requirements, allowing us to evaluate our method’s scalability and generalization across different task difficulties. While Robotouille represents a specific domain, its structured nature with consistent state-action distribution under varying complexity levels makes it particularly suitable as a controlled yet comprehensive testbed to empirically validate our theoretical results. Future work will extend it to more diverse agent environments [24, 25, 15] to demonstrate the broader applicability of single-turn decomposition for multi-turn agent training.
@@
Results of Cheese Sandwich SR ($\uparrow$) ASAT ($\downarrow$) ASST ($\downarrow$)
Llama3.3-70b (expert) 0.5 11.9 8.8
Qwen2.5-1.5b 0.0 15.0
Qwen2.5-3b 0.0 15.0
Qwen2.5-7b 0.0 15.0
Qwen2.5-14b 0.2 13.6 8.0
Qwen2.5-1.5b (w/ SFT) 0.0 15.0
Qwen2.5-1.5b (w/ SFT & GRPO) 0.3 12.5 6.7
@@
Llama3.3-70b (expert) 0.4 12.6 9.0
Qwen2.5-1.5b 0.0 15.0
Qwen2.5-3b 0.0 15.0
Qwen2.5-7b 0.1 14.9 14.0
Qwen2.5-14b 0.0 15.0 15.0
Qwen2.5-1.5b (w/ SFT) 0.2 13.7 8.5
Qwen2.5-1.5b (w/ SFT & GRPO) 0.7 10.1 8.0
@@
Llama3.3-70b (expert) 0.7 17.4 15.0
Qwen2.5-1.5b 0.0 23.0
Qwen2.5-3b 0.0 23.0
Qwen2.5-7b 0.0 23.0
Qwen2.5-14b 0.2 22.4 20.0
Qwen2.5-1.5b (w/ SFT) 0.6 18.4 15.3
Qwen2.5-1.5b (w/ SFT & GRPO) 0.7 15.8 12.7
@@
Llama3.3-70b (expert) 0.1 33.5 20.0
Qwen2.5-1.5b 0.0 35.0
Qwen2.5-3b 0.0 35.0
Qwen2.5-7b 0.0 35.0
Qwen2.5-14b 0.0 35.0
Qwen2.5-1.5b (w/ SFT) 0.1 34.2 27.0
Qwen2.5-1.5b (w/ SFT & GRPO) 0.3 30.5 20.0
Table 2: Cross-task Generalization of Success Rate (SR)
Evaluated ($\rightarrow$)
Trained ($\downarrow$)
Cheese
Sandwich
@@
Cheese
Burger
Double Cheese
Burger
Cheese Sandwich 0.3 0.3 0.0 0.0
Burger 0.4 0.7 0.0 0.0
Cheese Burger 0.4 0.7 0.7 0.0
Double Cheese Burger 0.3 0.5 0.4 0.3
Expert Trajectory Collection
Expert trajectories are collected using rejection sampling from Llama3.3-70B-Instruct [21], which serves as our expert policy $\pi^{GT}$. To ensure diversity and robustness, we generate expert trajectories across varied kitchen layouts and ingredient placements, creating approximately 100 successful trajectories per task type. Each trajectory $\tau^{GT}=[(s_{0}^{GT},a_{0}^{GT}),\ldots,(s_{n}^{GT},a_{n}^{GT})]$ satisfies the task completion criteria with $R(s_{n}^{GT},a_{n}^{GT})=1$. The state-action tokens follow the ReAct agentic framework [4], where each action $a_{i}^{GT}$ consists of structured reasoning followed by a specific action selection from the valid action list indicated from state tokens. This framework enables us to construct single-turn verifiable rewards $r(s_{i}^{GT},\pi_{\theta}(s_{i}^{GT}))=\mathds{1}\{\pi_{\theta}(s_{i}^{GT})=a_{i}^{GT}\}$ for GRPO training, where both the ReAct format and action correctness contribute to the reward by 0.1 and 0.9, respectively.
Model Training Pipeline
Our training pipeline consists of two sequential post-training stages using the VeRL implementation [26] of GRPO. First, we perform supervised fine-tuning (SFT) on Qwen2.5-1.5B-Instruct [22] using the collected expert trajectories to establish a strong initialization. The SFT stage trains the model to imitate the ReAct pattern for the action tokens. Subsequently, we apply GRPO with the single-turn verifiable reward function based on all state-action pairs $(s_{i}^{GT},a_{i}^{GT})$ from expert trajectories. The SFT takes 8 epochs with batch size of 8 and the GRPO training uses a batch size of 256, learning rate of $1\times 10^{-6}$, and KL regularization coefficient $\beta=0.001$ for 50 epochs.
Table 3: Cross-task Generalization of Average Steps of All Trajectories (ASAT)
Evaluated ($\rightarrow$)
Trained ($\downarrow$)
Cheese
Sandwich
Burger
Cheese
Burger
Double Cheese
Burger
Cheese Sandwich 12.5 13.1 23 35
Burger 13.5 10.1 23 35
Cheese Burger 12.2 10.8 15.8 35
Double Cheese Burger 13.3 11.6 19.4 30.5
Table 4: Cross-task Generalization of Average Steps of Successful Trajectories (ASST). “—” indicates no successful trajectories.
Evaluated ($\rightarrow$)
Trained ($\downarrow$)
Cheese
Sandwich
Burger
Cheese
Burger
Double Cheese
Burger
Cheese Sandwich 6.7 8.67
Burger 11.25 8.0
Cheese Burger 8.0 9.0 12.7
Double Cheese Burger 9.3 8.2 14 20
@@
We evaluate all models on 10 unseen kitchen layouts per task type, ensuring that the test environments differ from training data in terms of object positions, kitchen topology, and ingredient availability. The timeout lengths of attempt turns are 15, 15, 23, 35 for tasks of Cheese Sandwich, Burger, Cheese Burger, Double Cheese Burger, respectively. We assess the model performance of effectiveness and efficiency using three key metrics: (1) Success Rate (SR) - the proportion of episodes where the agent successfully completes the task within the step timeout limit; (2) Average Steps of All Trajectories (ASAT) - the mean number of steps taken across all episodes, including timeout failures; and (3) Average Steps of Successful Trajectories (ASST) - the mean number of steps for successfully completed episodes only. We compare our approach against different-size (1.5B, 3B, 7B, 14B) Qwen2.5 instruct models [22] using ReAct prompting [4] and the one right after SFT on expert trajectories but without GRPO training, ablating the contribution of RL for task reasoning.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in knowledge acquisition, reasoning, and tool use, making them promising candidates for autonomous agent applications. However, training LLM agents for complex multi-turn task planning faces significant challenges, including sparse episode-wise rewards, credit assignment across long horizons, and the computational overhead of reinforcement learning in multi-turn interaction settings. To this end, this paper introduces a novel approach that transforms multi-turn task planning into single-turn task reasoning problems, enabling efficient policy optimization through Group Relative Policy Optimization (GRPO) with dense and verifiable reward from expert trajectories. Our theoretical analysis shows that GRPO improvement on single-turn task reasoning results in a lower bound of the multi-turn success probability under the minimal turns, as well as the generalization to subtasks with shorter horizons. Experimental evaluation on the complex task planning benchmark demonstrates that our 1.5B parameter model trained with single-turn GRPO achieves superior performance compared to larger baseline models up to 14B parameters, with success rates of 70% for long-horizon planning tasks.
@@ 1 Introduction
  • We provide a theoretical analysis bridging single-turn task reasoning and multi-turn task completion, proving that the success probability after GRPO can be lower-bounded in multi-turn planning scenarios.
@@ 2.2 Expert Trajectory with Unique Optimality
On top of Definition˜2.1, we further define the expert trajectory generated by the expert policy with the unique optimality of minimal trajectory length. We adopt superscript GT{GT} to denote “Ground Truth” for expert policy or trajectory.
@@
Even though expert trajectory $\tau^{GT}(T^{GT})$ is uniquely optimal with minimal length of turns $T^{GT}$, the induced expert policy $\pi^{GT}$ is not necessarily unique or deterministic, as there is no state-action constraint beyond expert trajectories $\tau^{GT}(T^{GT})$. The unique optimal expert trajectory assumption ensures a well-defined single-step reward and serves as the foundation for the following results. We ensure this assumption in the experiment and further conduct an ablation comparison of robustness under non-expert noisy training trajectories.
@@ 3 Methodology
We first give the theoretical analysis of policy learning for single-turn MDP $\mathcal{M}_{S}$ based on expert trajectories.Then we define the success probability with the minimal turns for task planning MDP $\mathcal{M}$ and show that policy improvement from $\mathcal{M}_{S}$ can give a lower bound of success probability in $\mathcal{M}$, for even unseen task planning.
@@ 3.1.2 GRPO Single-Turn Optimality and Success Amplification
The optimal policy at iteration $n$ of the GRPO dynamics in Equation˜6 satisfies $\pi_{n}(a|s)\propto\pi_{\text{ref}}(a|s)\exp\left(\frac{1}{\beta}A_{\pi^{GT}}(s,a)\right),$ where the advantage function $A_{\pi^{GT}}$ upweights successful actions ($r_{\pi^{GT}}(s,a)=1$) and downweights failures ($r_{\pi^{GT}}(s,a)=0$). The variance normalization in $A_{\pi^{GT}}$ ensures positive advantages for successes and negative advantages for failures. At the fixed point $\pi^{*}$, this means $\exp(\frac{1}{\beta}A_{\pi^{GT}})>1$ for correct actions and $\exp(\frac{1}{\beta}A_{\pi^{GT}})<1$ for incorrect ones, systematically shifting probability mass toward successful outputs and yielding $p^{*}>p_{\text{ref}}$. Check Appendix C.1 of [19] for full proof. ∎
@@ 3.2 Success Probability for Multi-Turn Task Planning
Please note that the recursion equation in Theorem˜3.2 is not the same as Bellman recursion in value function iteration, and the key difference lies in that the multiplication of the binary single-turn reward $r_{\pi^{GT}}(s_{t},a)$ is involved in the recursion of expectation, instead of addition. This recursion bridges the final multi-turn success in Equation˜8 with single-turn reward improvements at each step in Equation˜7, which is crucial to the following analysis on how single-turn GRPO gives a lower bound of success probability of the multi-turn task planning.
3.3 Lower Bound of Multi-Turn Success Probability
We now analyze how GRPO improvements on single-turn task reasoning in $\mathcal{M}_{S}$ will give a lower bound of success probability in the multi-turn task planning MDP $\mathcal{M}$. First, we assume the bounded Total Variation (TV) distance of actions between the proximal states in the training and test data. Then we show the lower bound of single-turn GRPO performance as Lemma 3.1.
Assumption 3.1 (Bounded Policy with State Proximity).
For any execution state $s\in\mathcal{S}_{\text{exec}}$ encountered during policy deployment, there exists a training state $s^{\prime}\in\mathcal{S}_{\text{train}}$ with the same expert action $\pi^{GT}(s^{\prime})=\pi^{GT}(s)$ such that for any policy $\pi$, $\|\pi(\cdot|s)-\pi(\cdot|s^{\prime})\|_{\text{TV}}\leq\delta$, where $\delta>0$ is a bounded constant and $\|\cdot\|_{\text{TV}}$ denotes the TV distance.
Lemma 3.1 (Lower bound of GRPO policy deployment).
Under Assumption 3.1, for any execution state $s\in\mathcal{S}_{\text{exec}}$, $\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}[r_{\pi^{GT}}(s,a)]\geq\mathbb{E}_{a\sim\pi^{\text{ref}}(\cdot|s)}[r_{\pi^{GT}}(s,a)]-4\delta,$ where $\pi^{*}$ is the GRPO-optimized policy and $\pi^{\text{ref}}$ is the reference policy.
Proof.
[…]
$\left|\mathbb{E}_{a\sim\pi(\cdot|s)}[r_{\pi^{GT}}(s,a)]-\mathbb{E}_{a\sim\pi(\cdot|s^{\prime})}[r_{\pi^{GT}}(s^{\prime},a)]\right|$ (16)
$\leq\sum_{a\in\mathcal{A}}|\pi(a|s)-\pi(a|s^{\prime})|\cdot|r_{\pi^{GT}}(s,a)|$
$\leq\sum_{a\in\mathcal{A}}|\pi(a|s)-\pi(a|s^{\prime})|=2\|\pi(\cdot|s)-\pi(\cdot|s^{\prime})\|_{\text{TV}}\leq 2\delta.$
By Corollary 3.1, since $s^{\prime}\in\mathcal{S}_{\text{train}}$:
$\mathbb{E}_{a\sim\pi^{*}(\cdot|s^{\prime})}[r_{\pi^{GT}}(s^{\prime},a)]$ $\geq\mathbb{E}_{a\sim\pi^{\text{ref}}(\cdot|s^{\prime})}[r_{\pi^{GT}}(s^{\prime},a)].$ (17)
Applying (16) to $\pi^{\text{ref}},\pi^{*}$ and combining (17) will complete the proof. ∎
Finally, the step-dependent lower bound of success probability is shown below, which becomes tighter as the time step is closer to the minimal length of turns $T^{GT}$ of expert trajectories. It is consistent with the intuition that the task is more likely to complete from the middle step compared to initial step.
Theorem 3.3 (Lower bound of GRPO success probability).
@@ 3.3 Lower Bound of Multi-Turn Success Probability
$P_{t}^{\pi^{*}}(s_{t})\geq P_{t}^{\pi^{\text{ref}}}(s_{t})-(T^{GT}-t)\cdot 4\delta,$ (18)
@@
By the existence and uniqueness of the expert trajectory in Definition˜2.2, denote the expert trajectory starting from $s_{t}$ as $\tau^{GT}=[(s_{t},a_{t}),\dots,(s_{T^{GT}},a_{T^{GT}})]$ with minimal turns of task completion $T^{*}(s_{t})=T^{GT}-t$ in Definition˜3.1. We prove by backward induction on the length of remaining turns to task completion. In the final case when $t=T^{GT}$, we have $P_{t}^{\pi^{*}}(s_{t})=P_{t}^{\pi^{\text{ref}}}(s_{t})=1$ since the task is already completed. Now for the inductive case at step $t+1,t+1\leq T^{GT}$, we assume $P_{t+1}^{\pi^{*}}(s^{\prime})\geq P_{t+1}^{\pi^{\text{ref}}}(s^{\prime})-(T^{GT}-t-1)\cdot 4\delta$ for all $s^{\prime}\in\mathcal{S}$. Then for any state $s$ at step $t$, by Theorem˜3.2:
@@
$\geq\mathbb{E}_{a\sim\pi^{*}}[r_{\pi^{GT}}(s,a)\{P_{t+1}^{\pi^{\text{ref}}}(f(s,a))-(T^{GT}-t-1)4\delta\}]$ (20)
$=\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}[r_{\pi^{GT}}(s,a)\cdot P_{t+1}^{\pi^{\text{ref}}}(f(s,a))]-(T^{GT}-t-1)\cdot 4\delta\cdot\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}[r_{\pi^{GT}}(s,a)]$ (21)
$\geq\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}[r_{\pi^{GT}}(s,a)\cdot P_{t+1}^{\pi^{\text{ref}}}(f(s,a))]-(T^{GT}-t-1)\cdot 4\delta$ (22)
$\geq\mathbb{E}_{a\sim\pi^{\text{ref}}(\cdot|s)}[r_{\pi^{GT}}(s,a)\cdot P_{t+1}^{\pi^{\text{ref}}}(f(s,a))]-4\delta-(T^{GT}-t-1)\cdot 4\delta$ (23)
$=P_{t}^{\pi^{\text{ref}}}(s)-(T^{GT}-t)\cdot 4\delta,$ (24)
where (20) applies the inductive hypothesis, (22) uses $\mathbb{E}_{a\sim\pi^{*}(\cdot|s)}[r_{\pi^{GT}}(s,a)]\leq 1$, (23) applies Lemma 3.1 (noting that since $0\leq P_{t+1}^{\pi^{\text{ref}}}(s^{\prime})\leq 1$, the weighted expectation difference is bounded by the marginal difference), and (24) uses Theorem˜3.2 for $\pi^{\text{ref}}$. ∎
@@
For any subtask defined by completion step $k^{*}<T^{GT}$ in Definition˜3.3, the GRPO-optimized policy $\pi^{*}$ achieves better performance on the subtask than the reference policy $\pi^{\text{ref}}$ for any given state $s_{t}$ at step $t\leq k^{*}$, i.e., $P_{t}^{\text{sub},\pi^{*}}(s_{t})\geq P_{t}^{\text{sub},\pi^{\text{ref}}}(s_{t})-(k^{*}-t)\cdot 4\delta$.
@@ 4 Experiments
In the experiments, we aim to answer the following two questions empirically: 1) How is the performance of multi-turn task planning after GRPO over single-turn task reasoning? 2) Are GRPO-trained agents generalizable to unseen planning tasks and robust to noisy non-expert training trajectories? We answer the first question in Section˜4.2 and the second one in Section˜4.3. Prior to the answers, we first introduce the experimental setup of environment, model training and evaluation metrics.
@@ 4.1 Experimental Setup
We evaluate our approach on the Robotouille benchmark [20], a challenging multi-turn task planning environment designed for testing LLM agents’ reasoning capabilities in cooking scenarios. Robotouille provides a deterministic long-horizon complex environment where agents must coordinate multiple actions across different kitchen stations to complete recipes. We focus on three progressively challenging cooking tasks with different maximum steps: (1) Burger - at most 10 steps; (2) Cheese Burger - at most 15 steps; and (3) Double Cheese Burger - at most 23 steps. These tasks were selected to provide increasing complexity in both planning horizon and reasoning requirements while sharing the in-distribution states and actions, allowing us to evaluate our method’s scalability and generalization across different task difficulties.
@@
[…]
@@
[…]
@@
[…]
Table 2: Performance comparison on Burger task with different RL post-training methods. Best results are in bold.
RL post-training methods SR ($\uparrow$) ASAT ($\downarrow$) ASST ($\downarrow$)
[…]
Table 3: Cross-task generalization of SR and ASAT.
Evaluation Tasks ($\rightarrow$)
Model Training Tasks ($\downarrow$)
Burger SR /
ASAT
Cheese Burger
SR / ASAT
Double Cheese
Burger SR / ASAT
@@
[…]
$\pm0.0$ /
[…]
$\pm0.0$ /
[…]
Cheese Burger
[…]
$\pm0.0$ /
[…]
Double Cheese Burger
[…]
Table 4: Performance comparison on Burger task with different ratio of noisy trajectories. Best results are in bold.
Ratio of noisy trajectories SR ($\uparrow$) ASAT ($\downarrow$) ASST ($\downarrow$)
[…]
Data Collection and Model Training
Expert trajectories are collected using rejection sampling and minimum-step trajectory filtering from Llama3.3-70B-Instruct [21], which serves as our expert policy $\pi^{GT}$. To ensure diversity and robustness, we generate expert trajectories across varied kitchen layouts and ingredient placements, creating approximately 100 successful trajectories per task type. For the ablation of training on noisy non-expert trajectories, we randomly choose 10%, 30%, 50% of the states along all trajectories that are perturbed to have sub-optimal actions but keep the task completion successful. The state-action tokens follow the ReAct agentic framework [4], where each action $a_{i}^{GT}$ consists of structured reasoning followed by a specific action selection from the valid action list indicated from state tokens. ReAct format and single-turn correctness reward of Definition˜2.4 are weighted by 0.1 and 0.9, respectively. Our training pipeline consists of two sequential post-training stages using the VeRL implementation [22] of RL post-training. First, we perform supervised fine-tuning (SFT) on Qwen2.5-1.5B-Instruct [23] using the collected expert trajectories to establish a strong initialization. The SFT takes 8 epochs with batch size of 8 and the RL post-training uses a batch size of 256, learning rate of $1\times 10^{-6}$, and KL regularization coefficient $\beta=0.001$ for 50 epochs.
@@
We evaluate all models on 10 unseen kitchen layouts per task type, ensuring that the test environments differ from training data in terms of object positions, kitchen topology, and ingredient availability. The timeout lengths of attempt turns are 15, 23, 35 for tasks of Burger, Cheese Burger, Double Cheese Burger, respectively. We assess the model performance of effectiveness and efficiency using three key metrics: (1) Success Rate (SR) - the proportion of episodes where the agent successfully completes the task within the step timeout limit; (2) Average Steps of All Trajectories (ASAT) - the mean number of steps taken across all episodes, including timeout failures; and (3) Average Steps of Successful Trajectories (ASST) - the mean number of steps for successfully completed episodes only. The temperature is set to be 0.7 and all values are in Mean $\pm$ Standard Deviation from 3 random trials. We compare our approach against prompt-based larger agents (7B, 14B models of Qwen2.5), 1.5B models only with SFT, and 1.5B models with SFT and other RL post-training methods (PPO [24], RLOO [25], REINFORCE++ [26]).